Evaluate the following expression:
C(n,2)
I understand that it would be:
= n! / r!(n-r)!
= n! / 2!(n-2)!
I'm just not sure how to simplify this?
An explanation and answer would be really appreciated :)

Answer by ewatrrr(24785)

= n! / 2!(n-2)! = (n(n-1)(n-2)(n-3)...)/(2!(n-2)(n-3)...) = n(n-1)/2
Ex:
5C2 = 5*4/2 = 10